### v1

# Enabling local editing in Diffusion Models

[…]
Enabling local editing in Diffusion Models
Appendices
In this document, we present additional material to support the main paper. Firstly, in Section A for reproducibility purposes we provide implementation details and detailed algorithms for our method. Lastly, in Section B we provide additional experimental results, comparisons and an ablation study designed to supplement and further validate our proposed method.
A Algorithms & Implementation Details
A.1 Algorithms
In this Section, we provide an algorithm that describes our method as presented in Section 4.2 in the main paper. Additionally, we offer detailed algorithms for the Subspace Iteration method used to approximate the SVD of the Jacobian of the denoising network and the JIVE algorithm used to obtain a solution to the minimization problem in Eq. (5) of the main paper.
Computing Joint and Individual Components in the Latent Space of DMs: Algorithm 1 summarizes our method for decomposing the Jacobian of each region into a joint and an individual component. We start with a real image $\mathbf{x}_{0}$, a denoising network $\mathbf{e}_{\theta}$ and a set of regions $M=\{\mathbf{m}_{1},\cdots,\mathbf{m}_{N}\}$, along with the denoising timestep $t$ and joint and individual ranks $r_{C}$ and $r_{A}$. Starting with $\mathbf{x}_{0}$, we obtain its corresponding noise $\mathbf{x}_{T}$ via DDIM Inversion [song2020denoising]. We denoise $\mathbf{x}_{T}$ up to timestep $t$ with the standard DDIM reverse process (see Eq. (3) in text). Then we obtain the SVD of the Jacobian of each region with the Subspace Iteration (Algorithm 2). Finally, by utilizing the JIVE (Algorithm 3) we obtain the joint and individual components $\mathbf{C}$ and $\mathbf{A}_{i}$ for each region $i$. The rows of $\mathbf{A}_{i}$ are latent directions that result in meaningful local edits within region $i$.
Algorithm 1 Computing Joint and Individual Components in the Latent Space of DMs
1:procedure ComputeIndivJoint($\mathbf{e}_{\theta},\mathbf{x}_{0},t,\text{M},r_{C},r_{A}$)
2:     $\mathbf{x}_{T}\leftarrow\text{DDIMInversion}(\mathbf{x}_{0})$triangleright DDIM Inversion [song2020denoising]
3:     $\mathbf{x}_{t}\leftarrow\text{DDIMReverse}(\mathbf{x}_{T},t)$$\triangleright$ DDIM Reverse process until timestep t
4:     for $i,\mathbf{m}\;\;\text{in}\;\;\text{enumerate}(\text{M})$ do $\mathbf{U}^{(i)},\mathbf{S}^{(i)},\mathbf{V}^{(i)}\leftarrow\text{SubspaceIteration}\big{(}\{\mathbf{e}_{\theta}^{t}(\mathbf{x}_{t})\}_{m}\big{)}$triangleright Algorithm 2 $\mathbf{J}_{\perp}^{(i)}\leftarrow\mathbf{S}^{(i)}\mathbf{V}^{(i)^{T}}$triangleright Dimension-reducing transformation
5:     end for
6:     $\mathbf{J}_{\perp}=\big{[}\mathbf{J}^{(1)^{T}},\cdots,\mathbf{J}^{(N)^{T}}\big{]}^{T}$
7:     $\mathbf{C},\{\mathbf{A}_{i}\}_{i=1}^{N}\leftarrow\text{JIVE}(\mathbf{J}_{\perp},r_{C},r_{A})$triangleright Algorithm 3
8:     return $\mathbf{C},\{\mathbf{A}_{i}\}_{i=1}^{N}$
9:end procedure
Jacobian Subspace Iteration: As discussed in Section 4.2 the dimension of the latent space and the output image, result in a Jacobian matrix of approximately 6B parameters. To efficiently calculate the SVD of the Jacobian of the denoising network without storing it in memory we rely on Jacobian Subspace Iteration as proposed by Haas et al. [haas2023discovering]. Note that differently from [haas2023discovering] we calculate the SVD of the Jacobian of the denoising network explicitly within a region of interest $\mathbf{m}$.
Algorithm 2 Jacobian Subspace Iteration
1:procedure SubspaceIteration($\mathbf{e}_{\theta},\mathbf{x}_{t},\mathbf{m}$)
2:     $\mathbf{h}_{t}\leftarrow\mathbf{e}_{\theta}^{t}(\mathbf{x}_{t})_{:\frac{l}{2}}$$\triangleright$ Apply only the encoder to get bottleneck featuremaps
3:     $\mathbf{y}_{m}\leftarrow\{\mathbf{e}_{\theta}^{t}(\mathbf{x}_{t})\}_{m}$$\triangleright$ Output of denoising network in region $\mathbf{m}$
[…]
5:     $\mathbf{Q},\mathbf{R}\leftarrow\text{QR}(\mathbf{V})$triangleright Reduced QR decomposition
6:     $\mathbf{V}\leftarrow\mathbf{Q}$triangleright s.t. $\mathbf{V}\mathbf{V}^{T}=\mathbf{I}$
7:     while termination criterion do $\mathbf{U}\leftarrow\frac{\partial\{\mathbf{e}_{\theta}^{t}(\mathbf{x}_{t}|\ \mathbf{h}_{t}+a\mathbf{V})\}_{m}}{\partial a}\Big{|}_{a=0}$triangleright Forward differentiation $\hat{\mathbf{V}}\leftarrow\frac{\partial\mathbf{U}^{T}\mathbf{y}_{m}}{\partial\mathbf{h}_{t}}$ $\mathbf{V},\mathbf{S},\mathbf{R}\leftarrow\text{SVD}(\hat{\mathbf{V}})$triangleright Reduced SVD
8:     end while
9:     return $\mathbf{U},\mathbf{S},\mathbf{V}$,
10:end procedure
Joint and Individual Variation Explained (JIVE) Algorithm: For completeness we provide the iterative JIVE algorithm [lock2013joint]. Given a set of $N$ matrices $\mathbf{X}=[\mathbf{X}^{(1)}\cdots\mathbf{X}^{(N)}]$, JIVE iteratively approximates their joint and individual components.
Algorithm 3 Joint and Individual Variation Explained
1:procedure JIVE($\mathbf{X},r_{C},r_{A}$)
2:     $\mathbf{X}_{joint}\leftarrow\big{[}\mathbf{X}^{{(1)}^{T}}\cdots\mathbf{X}^{{(N)}^{T}}\big{]}^{T}$
3:     while termination criterion do
4:         $\mathbf{C}=\big{[}\mathbf{C}^{{(1)}^{T}},\cdots,\mathbf{C}^{{(N)}^{T}}\big{]}\leftarrow\text{rank}\;r_{C}\;\text{SVD of}\;\mathbf{X}_{joint}$
5:         for $i=(1,\cdots,N)$ do
6:              $\mathbf{X}^{(i)}_{indiv}\leftarrow\mathbf{X}^{(i)}-\mathbf{C}^{(i)}$
7:              $\mathbf{A}^{(i)}\leftarrow\;\;\text{rank}\;r_{A}\;\text{SVD of}\;\mathbf{X}^{(i)}_{indiv}(\mathbf{I}-\mathbf{V}\mathbf{V}^{T})$triangleright Ensures orthogonality constraint
8:              $\mathbf{X}^{(i)}_{joint}\leftarrow\mathbf{X}^{(i)}-\mathbf{A}^{(i)}$
9:         end for
10:         $\mathbf{X}_{joint}=\big{[}\mathbf{X}^{{(1)}^{T}}_{joint}\cdots\mathbf{X}^{{(N)}^{T}}_{joint}\big{]}^{T}$
11:     end while
12:     return $\mathbf{C},\{\mathbf{A}\}_{i=1}^{N}$
13:end procedure
A.2 Impementation Details
Here we describe more thoroughly the experimental setup we use to produce our results in the quantitative comparisons in the main paper, in order to ensure reproducibility. For all methods compared we use an editing strength $a=50$ along the identified semantic direction to obtain the edits. The alternative methods we benchmark our proposed method against are detailed below.
Asyrp For Asyrp [kwon2022diffusion] we use the author’s official code [^1] https://github.com/kwonminki/Asyrp_official. Since Asyrp is a supervised approach we train each attribute with 1k samples following the recipe detailed in their paper. Asyrp uses CLIP to obtain the editing directions and requires a source caption $y^{\text{source}}$ and a target caption $y^{\text{target}}$. The source and target captions we use to obtain the editing directions for each attribute are:
  • Smile: $y^{\text{source}}=$face", $y^{\text{target}}=$face with a smile"
  • Gaze: $y^{\text{source}}=$face", $y^{\text{target}}=$face looking left"
  • Red Lips: $y^{\text{source}}=$face", $y^{\text{target}}=$face with red lips"
  • Close Eyes: $y^{\text{source}}=$face", $y^{\text{target}}=$face with closed eyes"
Haas et al. We implement the method of Haas et al. [haas2023discovering] following their paper  [^2] At the time of writing the paper the official codebase of Haas et al. was not public yet.. The editing directions for Smile, Close Eyes, Red Lips, and Gaze edits were obtained from CelebA-HQ samples with index numbers 00009, 00000, 00042, 00003 respectively.
Ours For our method we obtain all editing directions from CelebA-HQ sample with index number 00020.
B Additional Experimental Resuls
B.1 Ablation Study: Choice of joint and individual ranks
Figure 1: Joint and individual ranks ablation. The editing result with joint and individual ranks used for all our experiments i.e. $r_{C}=30,\;r_{A}=5$ is highlighted with a green rectangle.
Here we turn our focus on the effect of joint and individual rank selection on local editing. In Figure 1 we present the editing results for a localized attribute manipulation (open eyes) under various joint and individual rank choices. We highlight with a green rectangle the editing result obtained $r_{C}=30$ and $r_{A}=5$ which are the ranks used for all experiments. We observe that a low joint rank results in undesirable edits outside the region of interest, like changing the mouth and the hairband. Similar effects are observed under a high individual rank. Conversely, a low individual rank and a high joint rank produce a localized manipulation that minimally affects the rest of the image.
B.2 Edits from different timesteps
Figure 2: Editing results from different timesteps of the denoising process for the CelebA-HQ datasets Regions of interest are denoted by a pink rectangle. All edits presented correspond to the $1^{\text{st}}$ principal component for each timestep.
In the main paper, we present editing results obtained at timestep $t=0.6\text{T}$ of the denoising process. In Figure 2 and Figure 3 we present editing results obtained from various timesteps on CelebA-HQ, LSUN-Churches, and MetFaces. All edits presented correspond to the first principal component for each timestep. We observe that our method is robust with respect to timestep selection, identifying meaningful semantic directions at various timesteps. For CelebA-HQ, these directions produce edits like changing the eye color or the gaze for the eyes region and changing the lip color and the expression for the mouth region. For LSUN-Churches, editing the window region produces different window variations, while for MetFaces, editing the mouth region changes the subject’s expression and facial hair.
Figure 3: Editing results from different timesteps of the denoising process for the LSUN-Churches and MetFaces datasets. Regions of interest are denoted by a pink rectangle. All edits presented correspond to the $1^{\text{st}}$ principal component for each timestep.
B.3 Linear Interpolation between directions
Figure 4: Linear Interpolation between directions for a sample from CelebA-HQ.
In Figure 4 we showcase how latent directions identified by our method can be composed to simultaneously edit different attributes. Let $\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3},\mathbf{v}_{4}$ be the latent directions corresponding to Gaze, Red Lips, Brows, and Smile edits respectively. Each image in position $(i,j)$ in Figure 4 is produced with editing vector $\mathbf{v}_{i,j}=a_{i}\sum_{k=1}^{j}\mathbf{v}_{k}$ with $a_{i}\in[-40,...,40]$ and $\mathbf{v}_{k}\in[\mathbf{v}_{1},\mathbf{v}_{2},\mathbf{v}_{3},\mathbf{v}_{4}]$, where $a_{i}$ is the editing strength. We observe that scaling the strength of the edit controls the magnitude of attribute change and that the negative scales result in semantically opposite edits.
B.4 Additional Comparisons
Figure 5: Additional qualitative comparisons between our method and existing alternatives for Red Lips and Gaze edits.
In Figure 5 we present additional qualitative comparisons for two semantic directions not shown in the main paper. Asyrp [kwon2022diffusion] produces considerable distortions and artifacts in the edited regions and in some instances significantly alters the subject’s identity, as in the second and third images of the Red Lips edit. Haas et al. [haas2023discovering] are better at preserving the subject’s identity but the edits fail to be localized to the region of interest, changing the subject’s overall expression. Our method succeeds in producing the desired semantic edits in the region of interest with almost imperceptible changes to the other parts of the image.